What is the equation of a line passing through (-6 , 5) and having a slope of 1/3
Answer by stanbon(75887)

What is the equation of a line passing through (-6 , 5)
and having a slope of 1/3
-------
Form: y = mx + b
Solve for "b":
5 = (1/3)(-6)+b
5 = -2 + b
b = 7
------
Equation:
y = (1/3)x + 7
==================
Cheers,
Stan H.
==================